Circular No. 8535

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 K3 (McNAUGHT) R. H. McNaught reports his discovery of a comet (discovery position

below) with the 0.5-m Uppsala Schmidt in the course of the Siding Spring Survey. On May 20.8 and 21.8 UT, there was a slightly diffuse head with a narrow tail 30′′ long in p.a. 240^{o}. On May 29.7, the appearance of the slightly diffuse head was much the same, but the tail was less obvious in the moonlight. On June 1.8, in excellent seeing, there was again a weak tail to the southwest, as well as an almost stellar central condensation.

2005 UT  α_{2000}  δ_{2000}  Mag.
May 20.79323  23^{h}42^{m}17.^{s}23  −11^{o}03′31′′.6  17.3

The initial astrometric observations, the following very preliminary elliptical orbital elements, and an ephemeris are given on MPEC 2005-L03.

T = 2005 Aug. 22.985 TT  ω = 24.^{o}063
e = 0.45244  Ω = 350.221 }2000.0
q = 1.43611 AU  i = 15.835
a = 2.62276 AU n^{o} = 0.232042 P = 4.25 years

COMET P/2005 L1 (McNAUGHT) McNaught reports his discovery of another comet (discovery position

below) with the 0.5-m Uppsala Schmidt in the course of the Siding Spring Survey. On June 2.7 UT, 60-s exposures showed a very weak, diffuse coma extending westward up to 10′′. After placement on the ‘NEO Confirmation Page’, observations were reported by others, including S. Foglia (Osservatorio Astronomico di Suno, 0.4-m f/4 reflector), who described the object as diffuse. On June 3.5, McNaught described the appearance as almost stellar, with a 10′′ diffuse tail to the west.

2005 UT  α_{2000}  δ_{2000}  Mag.
June 2.66008  16^{h}43^{m}25.^{s}63  −11^{o}51′02′′.4  16.7

The following orbital elements are from MPEC 2005-L30:

T = 2005 Dec. 15.369 TT  ω = 149.^{o}901
e = 0.20783  Ω = 138.300 }2000.0
q = 3.14774 AU  i = 7.745
a = 3.97358 AU n^{o} = 0.124432 P = 7.92 years
2005 June 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green